a ferry charges children $5 and adults $8. Let x be the number of children and y the number of adults. How many ordered pairs (combinations of cchildren and adults) consisting of whole numbers are possible so the total fee for al passengers is exactly $1000?
Answer by mananth(16949)

x+y=13

5x+8y=1000
8y=1000-5x
y=125-5/8 x
y has to be a whole number
so (5/8)x should be a whole number
x should take values of multipls of 8 till y is equal to 0
5/8 * x = 125
x=8*125/5 = 200

25 combinations are possible